# SimBEV: A Synthetic Multi-Task Multi-Sensor Driving Data Generation Tool and Dataset

### v1

@@ 3.4 Data annotations
SimBEV currently collects 3D object bounding boxes that are within a user-configurable radius of the ego vehicle (by default 120.0 m) at each time step for the following six classes: car, truck (includes trucks, vans, HGVs, etc., but not buses), bus, motorcycle, bicycle, and pedestrian. SimBEV collects other object attributes alongside each bounding box such as the object’s ID, its linear and angular velocities, and its make, model, and color if the object is a vehicle. An optional post-processing step calculates the number of lidar and radar points that fall within each bounding box and adds a valid label to boxes with at least one point inside, invalid otherwise. This labeling is useful when training 3D object detection algorithms, as it can filter out objects that are not guaranteed to be visible by the perception sensors [2].
BEV segmentation.
@@
Getting accurate BEV ground truth is the most critical aspect of SimBEV, which currently supports the following eight classes: road, car, truck, bus, motorcycle, bicycle, rider (human on a motorcycle or bicycle), and pedestrian. At each time step, the BEV ground truth is saved as a $C\times l\times l$ binary array, where $C$ is the number of classes and $l$ is the dimension of the BEV grid centered on the ego vehicle.
@@
To calculate the BEV ground truth for non-road classes, we leverage the fact that ground elements in CARLA (roads, sidewalks, etc.) use one-way visible materials, i.e. they are see-through from the bottom, as shown in Fig. 3. This means that we can place a semantic segmentation camera 1 km above the ego vehicle facing down (to minimize perspective distortion) and another one 1 km below the ego vehicle facing up to catch what the overhead camera may miss because of obstructions. Both cameras have a $l\times l$ resolution and the FoV is calculated such that each pixel represents a $d\times d$ area on the ground. The ground truth for each non-road class is obtained by merging data from the two cameras using a logical or operation. By default, $l$ is set to 360 and $d$ is set to 0.4 m, creating a 144 m $\times$ 144 m box around the ego vehicle. This is larger than what is typically used for BEV segmentation (100 m $\times$ 100 m), but the additional area can help with data augmentation (rotation, translation, scaling) during training.
Figure 3: Ground elements in CARLA use one-way visible materials.
@@
We follow an approach similar to [68] to get the ground truth for the road class. Specifically, we generate waypoints within a small distance of each other (specified by the user, we recommend setting it to $d$) across the map’s roads and note each waypoint’s lane width. We then calculate the mutual distance between these waypoints and the center of each cell of a $l\times l$ BEV grid centered on the ego vehicle where each cell represents a $d\times d$ area. For each grid cell, if a waypoint exists whose distance to the center of that cell is less than that waypoint’s lane width, that cell is labeled as road. Where our approach differs from [68] is that we then combine this information with data from the overhead camera and pass it through a binary closing operation to patch any potential holes, obtaining a more accurate ground truth.
@@
Overall, this approach results in accurate BEV ground truth data and also lets us assign multiple labels to the same cell. For instance, a cell occupied by a rider on a motorcycle will have a rider label (obtained from the overhead camera), a motorcycle label (obtained from the underground camera), and a road label (obtained using the waypoints).
@@
Our approach works everywhere except when roads with large elevation differences are present near the ego vehicle, e.g. when approaching an overpass. In those situations we do not use the semantic segmentation cameras. Instead, we only use the waypoint approach to calculate the road ground truth, and use 3D object bounding boxes to calculate the ground truth for other classes. While not as accurate as the previous method, the resulting ground truth is still acceptable. SimBEV intelligently switches between the two approaches when it detects two waypoints within 48.0 m of each other that have an elevation difference of at least 6.4 m, indicating they are on two different roads.
@@ 4 The SimBEV Dataset
To showcase SimBEV, we used it to generate the SimBEV dataset, a collection of 320 scenes spread across all 11 CARLA maps according to Tab. 2. Because Town12 and Town13 (CARLA’s two large maps) share many common features but use different building styles, textures, and vegetation, Town13 is not included in the train set to evaluate the generalization performance of trained perception models and expose overfitting issues.
@@
Table 2: Distribution of SimBEV dataset scenes across all available CARLA maps.
To create the SimBEV dataset, we used every sensor supported by SimBEV (RGB, semantic segmentation, instance segmentation, depth, and optical flow cameras; regular and semantic lidar; radar; GNSS; and IMU), positioned on the ego vehicle (a 2016 Ford Mustang) as shown in Fig. 4. We used a time step of 50 ms (i.e. 20 FPS), with each scene lasting 16 s (320 frames). In total, the SimBEV dataset contains 102,400 annotated frames, 8,315,935 3D bounding boxes (3,792,499 of which are valid), and 2,793,491,357 BEV ground truth labels. With 81.2 3D bounding boxes per frame (37.0 valid boxes per frame) it is on par with, if not surpassing, existing driving datasets [2, 4, 63, 62, 54]. A collection of front camera images displayed in Fig. 5 highlights the diversity of the SimBEV dataset.
@@
More information about the SimBEV dataset, including sensor properties, SimBEV parameters, and dataset statistics can be found in the Supplementary Material.
@@ 5 Evaluation and Analysis
The SimBEV dataset supports a variety of tasks including 2D and 3D segmentation, depth and optical flow estimation, tracking, prediction, and localization. Here, however, we focus on the two most common tasks: BEV segmentation and 3D object detection.
@@ 5.1 Tasks and metrics
We use the intersection over union (IoU) metric for the BEV segmentation task, where for each class, a prediction is considered positive if its probability (score) is above a certain threshold. Although we calculate the IoU for each class for thresholds from 0.1 to 0.9 in increments of 0.1, we only report the values for a 0.5 threshold here. Complete results can be found in the Supplementary Material.
Our 3D object detection task metrics are inspired by [2], and rely on the Average Precision (AP) metric. We use two different methods to define a match between a predicted 3D bounding box and a ground truth one. In the first, two boxes are considered a match when their 3D IoU is above a certain threshold [12, 6]. In the second, two boxes are considered a match if the distance between their centers is below a certain threshold. As explained in [2], small translation errors for small objects (such as pedestrians) result in a low or even zero IoU value, which makes comparing the performance of camera-only models that tend to have large localization errors difficult.
For both approaches, the AP is calculated as the area under the precision-recall curve. For the first method, we use matching IoU thresholds of $\mathbb{T}=\{0.3,0.4,0.5,0.6,0.7,0.8,0.9\}$. For the second method, similar to [2], we use matching distance thresholds of $\mathbb{T}=\{0.5,1,2,4\}$ m. For both approaches we define mAP as the average over all classes and all matching thresholds:
Similar to [2], we measure a set of True Positive metrics (TP metrics) for each predicted bounding box that is matched to a ground truth box. They are Average Translation Error (ATE), which is the Euclidean distance (in m) between box centers; Average Orientation Error (AOE), which is the smallest yaw angle difference (in rad) between the two boxes; Average Scale Error (ASE), which is equal to one minus the 3D IoU value of the two boxes after aligning for orientation and translation; and Average Velocity Error (AVE), which is the L2 norm of the difference in box velocities (in m/s). The mean TP metric (mTP) for each metric is computed by averaging over all classes and thresholds:
@@ 5.2 Evaluation results
We used BEVFusion [35], a state-of-the-art multi-sensor model for multi-task 3D perception, to evaluate both BEV segmentation and 3D object detection on the SimBEV dataset. BEVFusion offers camera-only, lidar-only, and fused (camera + lidar) model variants for each task (six variants in total), which lets us compare the performance of different sensing modalities. We trained all variants of BEVFusion using a batch size of 64, and sampled every 4th frame of the train set (17,600 samples in total). Furthermore, because of the imbalance in the number of objects for each class, we used the class-balanced grouping and sampling scheme utilized in the original implementation of BEVFusion [35, 76]. We used the validation set to tune a few hyperparameters, and evaluated the trained models on the test set. More information about BEVFusion and our training and evaluation strategy are provided in the Supplementary Material.
Modality Road Car Truck Bus Motorcycle Bicycle Rider Pedestrian Mean
C 74.1 26.7 17.0 29.5 0.0 0.0 0.0 0.0 18.4
L 89.6 75.8 77.5 84.8 44.6 4.0 38.5 27.7 55.3
C + L 89.9 77.0 79.3 84.8 45.7 4.5 37.7 32.7 56.5
Table 3: BEV segmentation IoU values (in %) for different variants of BEVFusion evaluated on the SimBEV test set.
Tab. 3 shows BEV segmentation IoU values (in %) for the three BEVFusion variants. As expected, all models achieve higher IoU values for larger objects compared to smaller ones (motorcycle, bicycle, rider, pedestrian). Moreover, IoU values for the road class (which is the only BEV segmentation class shared between SimBEV and nuScenes) are consistent with [35].
We can see from Tab. 3 that the fused model performs the best, followed closely by the lidar-only model because our lidar point cloud is very dense (approximately 225,000 points per frame). The fused model sees a 5% improvement in pedestrian IoU values, where it is helped by the additional semantic information from the camera. However, both models perform poorly when it comes to the bicycle class, though we found that the fused model achieves an IoU value of 17.9% for the bicycle class when the threshold is lowered to 0.4. It seems that because bicycles are always accompanied by a rider (and smaller and less noticeable than motorcycles), the model has difficulty distinguishing between the two and has lower confidence in its predictions.
Modality mAP mATE mAOE mASE mAVE SDS
@@
C 8.0 0.313 0.561 0.100 5.35 29.3
L 36.0 0.088 0.114 0.070 4.71 52.1
C + L 36.8 0.085 0.111 0.062 4.56 52.7
Table 4: 3D object detection results for different variants of BEVFusion evaluated on the SimBEV test set using the first method.
Tab. 3 also shows that the camera-only model performs poorly (except for the road class) compared to the other two, and reports a 0 IoU value for the smaller classes. As noted before, because images lack explicit geometric information, camera-only models have difficulty localizing objects.
3D object detection results using the first and second methods are shown in Tab. 4 and Tab. 5, respectively, and a comprehensive breakdown of the results is provided in the Supplementary Material. Like before, the fused model performs the best, followed closely by the lidar-only model. We can see that the second method produces higher mAP values, which is due to its more permissive nature where, unlike the first method, two boxes can be matched even if they do not intersect at all. This permissiveness, which makes the camera-only model more comparable to the other two, can be seen when juxtaposing the mATE, mAOE, and mASE values of the two methods, with those for the second method being considerably higher. Interestingly enough, this seems to indicate that both methods result in similar SDSs. As for the large mAVE values, because none of the input data has any embedded velocity information, it is hard for the model to estimate an object’s velocity. As noted in [2], aggregating consecutive point clouds decorated by their timestamp can significantly improve mAVE.
Modality mAP mATE mAOE mASE mAVE SDS
@@
C 22.7 0.708 0.686 0.131 4.92 29.8
L 46.2 0.151 0.201 0.109 4.62 54.9
C + L 46.3 0.148 0.198 0.097 4.48 55.1
Table 5: 3D object detection results for different variants of BEVFusion evaluated on the SimBEV test set using the second method.
Finally, we should note that the Swin-T [34] model used as BEVFusion’s image encoder was pretrained on ImageNet [8] and nuImages [2], which means the performance of models with image input may be less than optimal due to domain shift. Pretraining Swin-T using CARLA-based data may improve performance.
@@ 6 Conclusion
In this paper we introduce SimBEV, an extensively configurable and scalable randomized synthetic data generation tool that collects accurate BEV ground truth data and supports a variety of configurable perception and localization sensors. We hope that SimBEV enables users to rapidly create and scale different driving datasets, making it an invaluable tool for empowering researchers to explore various computer vision tasks. To showcase SimBEV, we use it to create the SimBEV dataset, a comprehensive large-scale driving dataset, which we use to benchmark BEV perception models and compare different sensing modalities and evaluation approaches. Future work will focus on expanding object classes supported by SimBEV, supporting custom sensor positions, and enabling V2X data collection.
@@ A CARLA Simulator
SimBEV relies on a modified version of CARLA Simulator 0.9.15 [9] with an enhanced content library. Some of the improvements we made are listed below.
  • We added three new sports cars to CARLA’s vehicle library using existing 3D models [22] [^1] We used royalty-free 3D models of the three cars available on BlenderKit as the basis for the vehicles. However, the Supra and Chiron models had been removed from BlenderKit at the time of writing, so unfortunately we have no way of crediting their authors for their work.: sixth generation Ford Mustang, Toyota GR Supra, and Bugatti Chiron, shown in Fig. 6. They increase the diversity of CARLA’s vehicle library, especially when it comes to fast, high-performance vehicles. The Ford Mustang is the default vehicle for collecting SimBEV data.
  • We added lights (headlights, taillights, blinkers, etc.) to some of the older vehicle models in CARLA’s library lacking them, and redesigned existing vehicle lights in Blender using a new multi-layer approach that better visualizes modern multi-purpose lights, as shown in Fig. 7.
  • We added a set of 160 standard colors for most vehicle models to choose from (apart from a few like the firetruck), and fixed color randomization issues for a few vehicles (e.g. the bus).
  • We updated the vehicle dynamics parameters of vehicle models to better match their vehicle’s behavior and performance in the real world.
@@
  • We updated the motorcycle and bicycle models so they select their driver models randomly each time, instead of always being assigned the same model.
@@
Figure 6: From left to right, the Bugatti Chiron, Ford Mustang, and Toyota GR Supra vehicles added to CARLA’s vehicle library with their lights turned off (top) and on (bottom).

### v2

@@ 3.4 Data annotation
At each time step, SimBEV collects 3D object bounding boxes that are within a user-configurable radius of the ego vehicle for the following six classes: car, truck (includes trucks, vans, HGVs, etc., but not buses), bus, motorcycle, bicycle, and pedestrian. Other object attributes are also collected alongside each bounding box, such as the object’s ID, its linear and angular velocity, and its make, model, and color if the object is a vehicle. An optional post-processing step calculates the number of lidar and radar points that fall within each bounding box and adds a valid label to boxes with at least one point inside, invalid otherwise. This labeling is useful for training 3D object detection algorithms, as it can filter out objects that may not be visible to perception sensors [3].
Figure 5: Left: BEV road data calculated using CARLA-generated waypoints; there are clear gaps where lanes diverge. Middle: BEV road data obtained from the overhead camera; vehicles and vegetation obstruct a portion of the view. Right: BEV road ground truth obtained by combining the two sources and performing binray closing.
BEV ground truth.
@@
SimBEV supports the following eight classes for BEV segmentation: road, car, truck, bus, motorcycle, bicycle, rider (human on a motorcycle or bicycle), and pedestrian. At each time step, the BEV ground truth is saved as a $C\times l\times l$ binary array, where $C$ is the number of classes and $l$ is the dimension of the BEV grid that is centered on the ego vehicle.
@@
To calculate the BEV ground truth for non-road classes, we take advantage of the fact that ground elements in CARLA (roads, sidewalks, etc.) use one-way visible materials, appearing solid from one direction and see-through from the opposite, as shown in Fig. 4. This means that we can place a semantic segmentation camera 1 km above the ego vehicle facing down (far enough to minimize perspective distortion) and another 1 km below the ego vehicle facing up to catch what the overhead camera cannot see due to obstructions. Both cameras have a $l\times l$ resolution and their field of view (FoV) is calculated so that each pixel represents a $d\times d$ area on the ground. The BEV ground truth for each non-road class is obtained by merging data from the two cameras using a logical or operation. By default, $l$ is set to 360 and $d$ is set to 0.4 m, creating a 144 m $\times$ 144 m box around the ego vehicle. This area is larger than what is typically used for BEV segmentation (100 m $\times$ 100 m), but it can help with data augmentation (rotation, translation, scaling) during training.
@@
We follow an approach similar to [64] to obtain the ground truth for the road class. Specifically, we use CARLA-generated waypoints a small distance apart from each other (specified by the user, we recommend setting it to $d$) across the map’s roads and note each waypoint’s lane width. We then calculate the mutual distance between these waypoints and the center of each cell of a $l\times l$ BEV grid that is centered on the ego vehicle, where each cell represents a $d\times d$ area. For each grid cell, if a waypoint exists whose distance to the center of that cell is less than that waypoint’s lane width, that cell is labeled as road. Where our approach differs from [64] is that we then combine this information with data from the overhead camera and perform binary closing to patch any potential gaps, obtaining a much more accurate ground truth. This process is illustrated in Fig. 5.
@@
In general, our method allows us to assign multiple labels to the same cell. For example, a cell occupied by a cyclist will have a rider (obtained from the overhead camera), a bicycle (obtained from the underground camera), and a road (calculated using CARLA-generated waypoints) label.
@@
Our approach works everywhere except when roads with large elevation differences are near the ego vehicle, e.g. when the ego vehicle is traveling under an overpass. In those situations, we do not use the overhead or underground cameras. Instead, we rely on CARLA-generated waypoints to calculate the BEV ground truth for the road class and use 3D object bounding boxes to calculate the BEV ground truth for other classes. Although not as accurate as our overall approach, the resulting ground truth is still acceptable. SimBEV switches to this method when it detects two waypoints within 48.0 m of each other that have an elevation difference greater than 6.4 m, signaling that they are on two different roads.
@@ 4 The SimBEV Dataset
To showcase SimBEV, we used it to create the SimBEV dataset, a collection of 320 scenes spread across all 11 CARLA maps according to Tab. 2. To the best of our knowledge, this is the first dataset that utilizes CARLA’s largest maps, i.e., Town12, Town13, and Town15. Because Town13 shares many common features with Town12 but uses different building styles, textures, and vegetation, it is not included in the train set to evaluate the generalization performance of trained models and expose overfitting.
@@
Table 2: Distribution of the scenes of the SimBEV dataset across all available CARLA maps.
@@
The SimBEV dataset was created on a computer with a single Nvidia GeForce RTX 3090 graphics card over the course of 83 hours. It takes up 1.5 TB when compressed, and contains data from every sensor supported by SimBEV. Data is collected at a 20 Hz sample rate, with each scene lasting 16 s (320 frames). In total, the SimBEV dataset contains 102,400 annotated frames, 8,315,935 3D bounding boxes (3,792,499 of which are valid), and 2,793,491,357 BEV ground truth labels. With 81.2 3D bounding boxes per frame (37.0 valid bounding boxes per frame) it is on par with, if not surpassing, existing driving datasets [3, 5, 59, 58, 50]. A collection of front camera images displayed in Fig. 6 highlights the diversity of the SimBEV dataset.
More information about the SimBEV dataset, including sensor properties, SimBEV parameters, and dataset statistics, can be found in the Supplementary Material.
@@ 5 Evaluation and Analysis
The SimBEV dataset can be used for a variety of perception tasks, including 2D/3D segmentation, depth and optical flow estimation, and motion tracking and prediction. Here, we focus on BEV segmentation and 3D object detection.
@@ 5.1 Tasks and metrics
BEV segmentation results are evaluated using intersection over union (IoU), where for each class, a prediction is considered positive if its probability (score) is above a certain threshold (here 0.5). Our 3D object detection metrics are inspired by [3], with average precision (AP) as the main metric. We consider two approaches to matching a predicted bounding box with a ground truth one. In the first, two boxes are matched if their 3D IoU is above a certain threshold [12, 7]. In the second, two boxes are matched if the distance between their centers is below a certain threshold. As [3] notes, in the former, small translation errors for small objects (such as pedestrians) result in low or even zero IoU, making performance comparison of camera-only models that tend to have large localization errors difficult. More information about the metrics used for 3D object detection evaluation is available in the Supplementary Material.
@@ 5.2 Evaluation results
Model Modality Road Car Truck Bus Motorcycle Bicycle Rider Pedestrian Mean
BEVFusion-C C 76.0 17.2 5.1 22.9 0.0 0.0 0.0 0.0 15.2
BEVFusion-L L 87.7 70.6 73.5 81.5 32.5 3.6 18.4 18.9 48.3
BEVFusion C + L 88.4 72.7 74.5 80.0 36.3 3.6 23.3 20.0 50.0
UniTR C + L 92.8 73.8 67.7 51.7 36.5 11.4 36.2 27.5 49.7
UniTR+LSS C + L 93.3 72.8 69.4 58.5 35.9 6.3 31.6 12.9 47.6
Table 3: BEV segmentation IoUs (in %) for different models evaluated on the SimBEV dataset test set. The top three values are indicated in red, green, and blue, respectively.
Model Modality mAP mATE mAOE mASE mAVE SDS
@@
BEVFusion-C C 7.0 0.337 0.943 0.106 4.98 23.7
BEVFusion-L L 33.9 0.105 0.086 0.107 1.49 50.8
BEVFusion C + L 34.1 0.107 0.077 0.101 1.46 51.0
UniTR C + L 33.0 0.081 0.140 0.071 0.51 56.5
UniTR+LSS C + L 34.2 0.083 0.131 0.069 0.49 57.5
Table 4: 3D object detection results for different models evaluated on the SimBEV dataset test set using the first (IoU-based) method. The top three values are indicated in red, green, and blue, respectively.
Model Modality mAP mATE mAOE mASE mAVE SDS
@@
BEVFusion-C C 22.1 0.744 1.044 0.137 4.65 25.1
BEVFusion-L L 48.1 0.144 0.133 0.134 1.56 56.4
BEVFusion C + L 48.1 0.146 0.122 0.127 1.54 56.6
UniTR C + L 47.7 0.113 0.224 0.090 0.55 61.7
UniTR+LSS C + L 47.8 0.113 0.207 0.085 0.53 62.2
Table 5: 3D object detection results for different models evaluated on the SimBEV dataset test set using the second (distance-based) method. The top three values are indicated in red, green, and blue, respectively.
We benchmark BEVFusion [33] and UniTR [53] - both multi-sensor models for multi-task perception - on the SimBEV dataset. BEVFusion has camera-only (BEVFusion-C), lidar-only (BEVFusion-L), and fused (camera + lidar) variants for each task (six variants in total), allowing us to compare the performance of different modalities. BEVFusion-C is a variant of BEVDet-Tiny [17] using a much heavier view transformer, while BEVFusion-L is the lidar-only variant of TransFusion (TransFusion-L) [1]. UniTR, along with the base model for each task, has a variant augmented by an additional LSS-based BEV fusion step (four variants in total) [33, 28, 43].
Table 3 shows BEV segmentation IoUs (in %) for different models evaluated on the SimBEV dataset test set. As expected, all models achieve higher IoUs for larger objects compared to smaller ones (motorcycle, bicycle, rider, and pedestrian). In addition, the IoUs for the road class (which is the only BEV segmentation class shared between SimBEV and nuScenes) are consistent with [33].
Table 3 shows that BEVFusion outperforms BEVFusion-L only by a small margin, probably because of SimBEV’s dense lidar point cloud. Notably, BEVFusion gets ahead when it comes to detecting smaller objects, probably because of the extra semantic information obtained from camera images. However, both models perform poorly when it comes to the bicycle class, though we found that BEVFusion achieves a 12.7% IoU for that class when the threshold is lowered to 0.4. It seems that because bicycles are always accompanied by a rider (and are smaller than motorcycles), the model has difficulty distinguishing between the two and has lower confidence in its predictions.
Table 3 also shows that BEVFusion outperforms UniTR, because the latter significantly underperforms in the bus class, even though it is much better at detecting smaller objects than the former. We think that this is likely because UniTR’s transformer backbone is unable to effectively utilize information in the $z$ direction. We can also see from Tab. 3 that BEVFusion-C performs poorly (except for the road class) compared to the others. As noted above, because images lack explicit geometric information, camera-only models have difficulty localizing objects.
3D object detection results using the first and second methods are shown in Tab. 4 and Tab. 5, respectively. In contrast to nuScenes benchmarks [3, 53], BEVFusion slightly outperforms UniTR and UniTR+LSS in mAP here. However, the UniTR variants score a much higher SDS because they do a much better job at predicting object velocities. We can also see that the second matching method (distance-based) produces higher mAP values. This is due to its more permissive nature, where, unlike the first matching method, two boxes can be matched even if they do not intersect at all. This permissiveness, which makes BEVFusion-C more comparable to the others, can be seen when juxtaposing the mATE, mAOE, and mASE values of the two methods, with those for the second method considerably higher. A breakdown of the results by class is available in the Supplementary Material.
Finally, we should note that while CARLA and the real world are statistically different domains, our results and those of [51] indicate that trends in CARLA are compatible with real-world observations and SimBEV, with its accurate BEV ground truth, can be a useful tool for evaluating both novel perception methods and domain adaptation strategies.
@@ 6 Conclusion
In this paper, we introduced SimBEV, a randomized synthetic data generation tool that is extensively configurable and scalable, supports a wide array of sensors, incorporates information from multiple sources to capture accurate BEV ground truth, and enables a variety of perception tasks including BEV segmentation and 3D object detection. To showcase SimBEV, we used it to create the SimBEV dataset, a comprehensive large-scale driving dataset, which we used to benchmark BEV perception models and compare different sensing modalities. We hope that SimBEV empowers researchers in exploring a variety of computer vision tasks. Future work will focus on improving SimBEV and enabling vehicle-to-everything (V2X) data collection.
@@ A CARLA Simulator
SimBEV relies on CARLA Simulator 0.9.15 [9] equipped with an enhanced content library. Some of the improvements we made are listed below.
  • We added three new sports cars to CARLA’s vehicle library using existing 3D models [22] [^1] We used royalty-free 3D models of the three cars available on BlenderKit as the basis for the vehicles. However, the Supra and Chiron models had been removed from BlenderKit at the time of writing, so unfortunately we have no way of crediting their authors for their work.: sixth generation Ford Mustang, Toyota GR Supra, and Bugatti Chiron, shown in Fig. 7. They enhance the diversity of CARLA’s vehicle library, especially when it comes to fast, high-performance cars. The Ford Mustang is the default data collection vehicle in SimBEV.
  • We added lights (headlights, taillights, blinkers, etc.) to some of the older models in CARLA’s vehicle library that lacked them, and redesigned existing vehicle lights in Blender using a new multi-layer approach that better visualizes modern multi-purpose lights, as shown in Fig. 8.
  • We added a set of 160 standard colors available to most vehicle models (apart from a few like the firetruck), and fixed color randomization issues for a few vehicles.
  • We updated vehicle dynamics parameters of vehicle models to better match their vehicle’s behavior and performance in the real world.
@@
  • We updated motorcycle and bicycle models so that they select their driver models randomly each time, instead of always being assigned the same model.
@@
Figure 7: From left to right, the Bugatti Chiron, Ford Mustang, and Toyota GR Supra added to CARLA’s vehicle library with their lights turned off (top) and on (bottom).